# Survival Multiarmed Bandits with Bootstrapping Methods

### v1

@@ Abstract
The Multiarmed Bandits (MAB) problem has been extensively studied and has seen many practical applications in a variety of fields. The Survival Multiarmed Bandits (S-MAB) open problem is an extension which constrains an agent to a budget that is directly related to observed rewards. As budget depletion leads to ruin, an agent’s objective is to both maximize expected cumulative rewards and minimize the probability of ruin. This paper presents a framework that addresses such dual goal using an objective function balanced by a ruin aversion component. Action values are estimated through a novel approach which consists of bootstrapping samples from previously observed rewards. In numerical experiments, the policies we present outperform benchmarks from the literature.
@@ 2 Ruin-Averse Bandits
where $\lambda$ is the hyper-parameter that drives the intensity of ruin aversion of the agent. [^1] Using the method of Lagrange multipliers, such problem would equivalent to solving $\max\mathbb{E}\left[b_{T}-b_{0}\right]$ subject to $\mathbb{P}[\tau\leq T]<c$ for some constant $c$ if arms distributions were known.
@@ 2.1 Action Value Definition
As such the greedy action is that which is estimated to maximize $\mathcal{O}_{t}(\lambda)$ if sampled repeatedly until $T$. Denote the time-$t$ action value of arm $k$ as
@@ 2.3 An Upper Confidence Bound Algorithm
Algorithms we propose to select respective actions is inspired from the Upper Confidence Bound (UCB) approach (see Auer et al., 2002). Such scheme tackles the exploration-exploitation dilemma by incorporating a term to the objective function that encourages the selection of actions that have not been sampled extensively previously and for which statistical confidence about their value is low. This promotes the idea of optimism in the face of uncertainty, which entail exploring actions based not only on their past performance, but also on their potential to reveal themselves as optimal once statistical uncertainty is resolved.

### v2

@@ Abstract
The Multiarmed Bandits (MAB) problem has been extensively studied and has seen many practical applications in a variety of fields. The Survival Multiarmed Bandits (S-MAB) open problem is an extension which constrains an agent to a budget that is directly related to observed rewards. As budget depletion leads to ruin, an agent’s objective is to both maximize expected cumulative rewards and minimize the probability of ruin. This paper presents a framework that addresses such a dual goal using an objective function balanced by a ruin aversion component. Action values are estimated through a novel approach which consists of bootstrapping samples from previously observed rewards. In numerical experiments, the policies we present outperform benchmarks from the literature.
@@ 2 Ruin-Averse Bandits
where $\lambda$ is the hyper-parameter that drives the intensity of ruin aversion of the agent. [^1] Using the method of Lagrange multipliers, such problem would be equivalent to solving $\max\mathbb{E}\left[b_{T}-b_{0}\right]$ subject to $\mathbb{P}[\tau\leq T]<c$ for some constant $c$ if arms distributions were known.
@@ 2.1 Action Value Definition
As such the greedy action is that which is estimated to maximize $\mathcal{O}_{t}(\lambda)$ if selected repeatedly until $T$. Denote the time-$t$ action value of arm $k$ as
@@ 2.3 An Upper Confidence Bound Algorithm
Algorithms we propose to select respective actions is inspired from the Upper Confidence Bound (UCB) approach (see Auer et al., 2002). Such scheme tackles the exploration-exploitation dilemma by incorporating a term to the objective function that encourages the selection of actions that have not been sampled extensively and for which statistical confidence about their value is low. This promotes the idea of optimism in the face of uncertainty, which entail exploring actions based not only on their past performance, but also on their potential to reveal themselves as optimal once statistical uncertainty is resolved.